OpenStudy (anonymous):

for odd "N" Prove that 3^n +7^n is divisible by 10

OpenStudy (anonymous):

if n is odd, then it can be factored (3+7)(3^(n-1) - 7*3^(n-2) ... + 7^(n-1))

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Do you know that a^n + b^n = (a+b) (a^(n-1) - ba^(n-2) ... + b^(n-1)

OpenStudy (anonymous):

nope, wait where is this rule from, cause the question i am doing atm is an induction question and i haven't seen such a rule

OpenStudy (anonymous):

then do you know that a^n - b^n = (a-b)(a^(n-1)... +b^(n-1)

OpenStudy (anonymous):

are you sure this works ....?

OpenStudy (mathmate):

a^n + b^n = (a+b) (a^(n-1) - ba^(n-2) ... + b^(n-1)) works if n is odd.

OpenStudy (mathmate):

Virtus, are you working on congruences?